OpenStudy (anonymous):

Give an algebraic proof that: 2n choose 2 = 2*(n choose 2) + n^2

OpenStudy (anonymous):

i think all you need is to write out what both sides actually are

OpenStudy (anonymous):

on the other hand, being algebra challenged today, either it is wrong or i am missing something

OpenStudy (anonymous):

that's right, just not sure about the expansion

OpenStudy (anonymous):

lhs is \(2n(2n-1)=4n^2-2n\) rhs is \(2n(n-1)+n^2=2n^2-2n+n^2=3n^2-2n\) so either they are not the same or i a need to go get more coffee

OpenStudy (zarkon):

you didn't divide by 2

OpenStudy (anonymous):

ho ho ho

OpenStudy (anonymous):

guess that would help wouldn't it?

OpenStudy (zarkon):

you do need to hit the coffee :)

OpenStudy (anonymous):

yeah i am really messing up today i blame the heat which is in fact why i am indoors on a saturday

OpenStudy (anonymous):

\(\binom{2n}{2}=\frac{2n(2n-1)}{2}=n(2n-1)\) \(2\binom{n}{2}+n^2=\frac{2n(n-1)}{2}+n^2\) now algebra